Brad's BBQ reported sales of $871,500 and net income of $43,000. Brad’s also reported ending total assets of $506,000 and beginning total assets of $399,000.

Required:

Calculate the return on assets, the profit margin, and the asset turnover ratio for Brad's BBQ.

RETURN ON ASSETS

CHOOSE NUMERATOR + CHOOSE DENOMINATOR = RETURN ON ASSETS

___________________+_______________________=_____________________

___________________+_______________________=_____________________

PROFIT MARGIN

CHOOSE NUMERATOR + CHOOSE DENOMINATOR = RETURN ON ASSETS

___________________+_______________________=_____________________

___________________+_______________________=_____________________
Answer #1

1) Return of Assets = net income / average total assets

= 43,000 / 452,500

= 9.50%

Average total assets = assets at beginning + assets at ending / 2

= 399,000 + 506,000 / 2

= 452,500

2) profit margin = net income / sale x 100

= 43,000 / 871,500 x 100

= 4.93%

3) assets turnover ratio = net sale / average total assets

= 871,500 / 452,500

= 1.93 times